# Sensitive parameter dependence of the self-similar dynamics in the bubble breakup in a confined geometry

### v5

@@ Abstract
Formation of a fluid drop, such as observed in dripping faucet, has been extensively studied for understanding singular dynamics widely observed in nature. The singular dynamics is often self-similar, i.e., the shape at different times are collapsed onto a master curve after rescaling. Here, we investigate the bubble breakup in an original setup, where a metal disk entrains air into viscous fluid with lubricating films on its surfaces in a confined cell, focusing on the dependence of the dynamics on the film thickness $e$. As a result, we find out the dynamics is self-similar as expected, but the shape of the master curve and the scaling exponents characterizing the self-similar dynamics are revealed to be sensitively dependent on $e$. We discuss analogy between the present study and critical phenomena in thermodynamic transitions, to suggest a promising future direction of the singular dynamics.
@@ I Introduction
Formation of a fluid droplet we observe on a daily basis, e.g., in a dripping faucet, has been extensively studied as representatives for singular dynamics widely observed in nature [1, 2], from gravitational collapse of stars [3] to cell aggregation [4]. Near the singularity, the shapes after rescaling often become independent from time collapsing on a master curve, i.e., the dynamics becomes self-similar [5]. Examples of the self-similar dynamics found in the formation of a fluid drop include breakup of a liquid drop in air, where inertia, capillary, and viscosity compete among others [6, 7], breakup of a bubble when high viscosity of the surrounding fluid competes with capillarity [8], and breakup of a bubble in a tube [9]. In all these examples, the self-similar dynamics of the breakup of a fluid drop is axisymmetric.
Recently, however, an example of the self-similar drop breakup without axisymmetry was reported [10] in a confined geometry and the breakup was further revealed to be suppressed if we further confined the system [11]. These unexpected findings were observed in an original system, in which a metal disk of radius $R$ ($\simeq$ cm) and of thickness $D_{0}$ ($\simeq$ mm) drags air into a viscous liquid phase with lubricating film of thickness $e$ ($\simeq$ mm) on the disk surfaces in a cell of thickness $D_{0}+2e$. However, in these previous studies, the shape dynamics was investigated mainly at $e=0.5$ mm. Here, we investigate the dependence of the shape dynamics on the film thickness $e$ for a fixed $D_{0}$ to find out that the shape dynamics changes sensitively with $e$ and the scaling exponents characterizing the self-similar dynamics are systematically dependent on $e$. This is in contrast with the previous studies [10, 11] in that the corresponding scaling exponents were found to be always the same value even when $D_{0}$ and $R$ were changed. We stress analogy between the self-similar dynamics and the scaling hypothesis in critical phenomena in thermodynamic transitions [12, 13]. This analogy suggests a promising avenue for the future study of singular dynamics by exploring symmetry in confined geometry.
@@
Figure 1: (a) Experimental setup. A metal disk of thickness $D_{0}$ ($=1$ mm) and radius $R$ (101010~{}10to $12.5$ mm) falls in the cell of thickness $D$ ($2$ to $6$ mm) filled with a viscous liquid of kinematic viscosity $\nu$ (100 to 1000 cS). The disk entrains air into the liquid, which finally detaches from the disk. The difference between $D$ and $D_{0}$ defines the liquid film thickness $e$. (b) Snapshots just before and at breakup illustrating the setting of axes, in the case with topology change for $(R,D_{0},e,\nu)=(10,3,0.5,100)$ in mm or cS.
@@
Figure 2: Snapshots of entrainment of air by a disk into liquid, leading to detachment of air from the disk for $R=10$ mm, $D_{0}=1$ mm, $\nu=100$ cS. The liquid film thickness $e$ are 1, 1.5, and 2 mm, respectively, in (a) to (c). The time label 0 ms corresponds to $t=t_{d}$ defined in the text. In the three rightmost front-view photos reveal that a small bubble remains on the solid surface in (b) and (c) while such a bubble cannot be seen in (a): ”the critical breakup” lies between $e=$ $1$ and 1.5 mm. Side-view shots near the detachment (0 ms) are separated by 6 ms, but are not synchronized with the front-view shots (a set of side-view snapshots are obtained from an experiment performed on a day different from the day on which the corresponding front-view snapshots but conducted for the same parameters).
@@ II.1 Experimental
In Fig. 1 (a), we show our experimental setup with explanation. The ranges of characteristic lengths, the radius and thickness of disk and the cell thickness, are as follows: $R=10-12.5$ mm, $D_{0}=1$ mm, $D=2-6$ mm. The cell width and height are much larger than the length scales $R,D_{0}$, and $D$ (typically 9 and 12 cm, respectively). We use polydimethylsiloxane (PDMS) for viscous liquid, where the range of kinematic viscosity $\nu=\eta/\rho$ is 100-1000 cS. The density $\rho$ and the surface tension $\gamma$ are slightly depending on viscosity $\eta$ ($\rho\simeq$ $0.97$ g/cm3 and $\gamma\simeq 20$ mN/m). The density $\rho_{s}$ of the metal disk (SUS430) is 7.7 g/cm3 with the density difference $\Delta\rho=\rho_{s}-\rho$. The cell is fabricated with acrylic plates of thickness 5 mm, using acrylic spacers whose thickness defines the cell thickness $D$.
@@ II.2 Shape of the air-liquid interface $h(z,t)$
In Fig. 1 (b), we explain the setting of coordinates in the present study (we set $z_{c}=0$, so that $z-z_{c}=z$ and $z_{m}-z_{c}=z_{m}$). The air-liquid interface formed by air entrained by the disk is seen as the dark curve with a finite width. The width reflects the three-dimensional characters of the interface: the interface is concave or convex. The region where the width is thin, the interface is almost flat. We track the right and left inner edges of the thick curve and describe them by $x=h(z,t)$ and $-h(z,t)$, respectively, except for the case of $e=0.5$ mm, for which we track instead the outer edges. We determine whether we track the inner or outer edge such that we can have a better collapse of shapes at different time after rescaling, which we discuss below. We call the minimum of the function $h(z,t)$ with respect to $z$ ”the constriction point” at which $(x,z)=(h_{m}(t),z_{m}(t))$, i.e., $h_{m}(t)=h(z_{m}(t),t)$.
II.3 Dynamics sensitive to the film thickness $e$
In Fig. 2, we show snapshots of air-detachment transitions near the air-solid contact for $e=1.0$, 1.5, and 2 mm, with the other parameters, $R$, $D_{0}$, and $\nu$ fixed (we have tried to analyze the case of $e=2.5$ mm, only to fail because the inner edge was not clear). In three rightmost front-view photos (labeled as 6 ms), detachment without topological change is observed for $e=1$ mm but that with topological change (i.e., breakup) for $e=1.5$ and $2$ mm: a small bubble remains on the solid surface in (b) and (c) while such a bubble cannot be seen in (a). In other words, ”the critical breakup” lies between $e=$ $1$ and 1.5 mm.
@@ II.3 Dynamics sensitive to the film thickness $e$
In the previous studies [10] and [11], the cases of $D_{0}\gtrsim\kappa^{-1}$ and $D_{0}\lesssim\kappa^{-1}$ with $\kappa^{-1}=\sqrt{\gamma/(\rho g)}\simeq 1.8$ mm are distinguished by the terms sheet-forming and corn-forming detachments, and the former and latter cases are respectively regarded as the cases with and without topology transition and with and without appearance of the constriction point. However, a careful examination on the $e$-dependence in the present study reveals that the previous distinction is applicable only for limited cases: in the present case, we observed both the cases with and without topology transition for $D_{0}\lesssim\kappa^{-1}$ and detachments shown in Fig. 2 are difficult to categorize by the sheet-forming or corn-forming case because it looks like the corn-forming case but the constriction point does appear.
@@ II.4 Dynamics of characteristic length scales
In Fig. 3 (a1) to (a4), we show the dynamics of three characteristic length scales, $2h_{m}(t)$, $z_{m}(t)$, and $z_{G}(t)$, for different film thickness $e$. The quantity $z_{G}(t)$ is the position of the center of gravity of the disk measured from that at the critical time $t=t_{c}$, which slightly differs from $t_{d}$ [$z_{G}(t_{c})=0$ by definition]. The critical time $t_{c}$ is determined by extrapolation as the time when $h_{m}$ becomes zero. The critical space-time point is represented as $(x,z,t)=(x_{c},z_{c},t_{c})$ with $x_{c}=h_{m}(t_{c})$ and $z_{c}=z_{m}(t_{c})$.
@@
$\begin{array}[c]{|c||c||c||c||c|}\hline\cr e\text{ (mm)}&\beta&\Delta&\delta&\Delta/(\beta\delta)\\ \hline\cr\hline\cr 0.5&1.0\pm 0.0&0.97\pm 0.06&1.01\pm 0.01&0.97\pm 0.06\\ \hline\cr 1.0&1.0\pm 0.0&0.79\pm 0.02&0.64\pm 0.00&1.2\pm 0.03\\ \hline\cr 1.5&0.93\pm 0.00&0.78\pm 0.01&0.61\pm 0.00&1.4\pm 0.02\\ \hline\cr 2.0&0.90\pm 0.01&0.77\pm 0.03&0.59\pm 0.00&1.5\pm 0.06\\ \hline\cr\end{array}$
We carefully selected the range of fitting indicated by a pair of cross marks for $h_{m}$ and $z_{m}$ to avoid including the data with large error bars and to use only the data with small error bars, by noting that (i) the error coming from the time resolution tends to become large near $t=t_{c}$ and (ii) the error tends to be large in $z_{m}$ compared with $h_{m}$. These points are visualized by the error bars shown for the data at $e=1.0$ mm in (b2) and (b3): The error bars in the horizontal axis are based on the time resolution ±$\pm 0.5$ ms at 2000 fps; The error bars in the vertical axis are on the spacial resolution corresponding to 1 pixel of the images for $h_{m}$ but possibly corresponding to several pixels for $z_{m}$ because of the elongation explained below. In accordance with Remark (i) and (ii), we can confirm the following in (b2) and (b3): the error bars in the horizontal axis is visible near $t=t_{c}$, while the error bars in the vertical axis are not visible for $h_{m}$ but significant for small $z_{m}$. The reason for large errors in $z_{m}$ is as follows. The constriction region tends to elongate near the breakup and the thin thread of air thus formed detaches from the disk when $h_{m}$ is not zero (see Fig. 6 below). Because of this elongation of the constriction region the determination of $z_{m}(t)$ tends to become difficult near breakup, which is reflected in larger standard deviation for $\Delta$ (compared with $\beta$) in Tab. 1.
As for the fitting of $z_{G}$, we justify the above linear fitting by Fig. 3 (c), which supports the following scaling law for $z_{G}(t)$ in agreement with the result established in a similar parameter range [11]:
@@ II.5 Self-similarity in the interface shape dynamics
In Fig. 4 (a1), we show the interfacial shape before rescaling at different times for $e=1$ mm. Interface shapes after rescaling are reasonably collapsed well on a master curve, especially near the constriction point, as seen in (a2). The collapse is significant for shapes before $t-t_{d}=-15$ ms as shown in (a3).
Figure 4: Shape change for $e=1$, 1.5, 2.0, and 0.5 mm for $(R,D_{0},\nu)=(10,1,100)$ in mm or cS. Since the right and left interfaces are the mirror image of the other, only the right interface after averaging is shown. The error bars in capturing the interface are less than the size of markers. The time development of interfacial shape before [after] rescaling is respectively shown in (a1) [(a2) and (a3)] for $e=1$ mm. We see a better collapse for times not too close to $t=t_{c}$ (before $-9$ ms) [see (a3)]. The corresponding plots for $e=$1.5, 2.0, and 0.5 mm are shown respectively in (b) to (d). The data labeled $R^{\ast}$ in (a3) are obtained for different parameter set $(R,D_{0},\nu)=(12.5,1,100)$, which collapse well on the master curve [the good collapse seems limited for the upper branch (of our focus), i.e., the shape above the constriction point].
Similar collapses are shown in (b) and (c) for $e=1.5$ and 2.0 mm, respectively. As $e$ increases the constriction region tends to elongate near breakup as announced, which can be confirmed by comparing (b1) with (c1) (see also Fig. 6 below). This elongation is related to the above-mentioned criticality of breakup: topology does change for $e=1.5$ and 2 mm with accompanied by breakup but it does not for $e=1.0$ mm. It seems that topological change (i.e., breakup) tends to elongate the constriction region near $t=t_{d}$. Because of this elongation, duration of a good collapse tends to be limited as $e$ increases [see (a2) to (c3) and (a3) to (c3)].
Reproducibility and universality are also demonstrated in Fig. 4. The data labeled $R^{\ast}$ in (a3) imply that the corresponding data are obtained for different $R$ with the other parameters fixed. In Fig. 4 (a3), even these data obtained for different $R$ collapse well on the master curve, which demonstrates the master curve is independent from (i.e., universal for a change in) $R$ for a fixed $e$, indirectly representing a good reproducibility of the experiment.
In Fig. 4 (d1), we show the interfacial shape for $e=0.5$ mm, which is fundamentally different from cases of $e$ larger than 0.5 mm. In this case, as mentioned above, we need to track the outer edge instead of inner edge of the dark interface to observe collapse of the renormalized shape as seen in (d2) [and more clearly in (d3)]. In addition, the constriction point does not appear [the point $(h_{m},$ $z_{m})$ is defined as the right contact point (at $y=0$) of the interface with the disk] but the tip forms a corn. However, the corn may not be axisymmetric but elliptic as reported in our previous study [11].

### v6

# Breakup to non-breakup transition of air entrained into viscous liquid by a disk: analogy of the self-similar dynamics with critical phenomena

@@ Abstract
Self-similarity in partial differential equations has been widely exploited to study many phenomena in physical sciences. We have studied the interfacial dynamics when air is entrained into viscous liquid by a disk in a confined geometry. In a previous study using an original experimental system, we found the sheet- and corn-forming regimes, in which a sheet and cone of air are respectively formed before air detaches from the disk. The sheet eventually breaks up but the corn, which appears when a bit more confined, does not. Here, we find a third regime, in which a corn eventually breaks up, by investigating different ranges of confining parameters: the transition from breakup to non-breakup can occur within the corn regime. Furthermore, with the data obtained in the third regime we deeply explore analogy with critical phenomena to find out that the counterpart of the critical exponents dependent on a length scale. Since the scale is a number not discrete but continuous, the present hydrodynamic analog suggests the existence of an uncountably infinite number of universality classes. The rich physics revealed in our study suggests a promising direction of the study of the self-similar dynamics: exploring analogy with critical phenomena, focusing on confined geometries in many natural and industrial phenomena.
@@ I Introduction
When shapes as a function of $x$ defined by a solution $h(t,x)$ of a partial differential equations (PDE) at different $t$ collapse onto a single master curve, the dynamics is called self-similar. The self-similarity in PDEs has played a vital role in understanding many phenomena in physical sciences [1, 2], in particular, phenomena categorized as the singular dynamics [3, 4], which include fluid-jet formation [5, 6], drop coalescence [7, 8, 9], gravitational collapse of stars [10] and cell aggregation [11]. Among them, the formation of a fluid droplet occurring, for example, in a dripping faucet, has been extensively studied. For the breakup of a liquid drop in air, a dynamical regime, in which inertia, capillary, and viscosity compete among others, is known [12, 13], while, for the breakup of a bubble, a regime, in which high viscosity of the surrounding fluid competes with capillarity, is found [14]. The latter regime is further found to exhibit a crossover to another viscous-capillary regime in the breakup of a bubble in a tube [15]. In all these examples, the self-similar dynamics of the breakup of a fluid drop is axisymmetric.
Recently, however, an example of the self-similar drop breakup without axisymmetry was reported [16] in a confined geometry. In the experiment, we developed an original experimental system, in which a metal disk entrains air into viscous liquid in a confined geometry, to find the sheet-forming regime with breakup, in which the entrained air eventually detaches from the disk with forming a bubble: before the detachment a non-axisymmetric constriction region appears in the interface where a sheet of air is formed, which thins down to breakup. In the second study, by using the same original experimental system, we further revealed another regime, the corn-forming regime without breakup. In this regime, a bubble is not created: the constriction point does not appear in the interface, but a corn of air, which is actually non-axisymmetric, is formed instead of a sheet near the detachment point, as a result of further confining the system [17]. (We remark here that these regimes were found in the pre-detachment dynamics, while the post-breakup dynamics in the sheet-forming regime has recently been studied [18].) Here, we find a third regime, the corn-forming regime with breakup, in the pre-detachment dynamics, in which the constriction region appears and then a bubble is created but a corn of air, instead of a sheet, is formed at the breaking tip, by investigating different ranges of confining parameters. In other words, we find that the transition from breakup to non-breakup can occur within the corn-forming regime, while in any cases the constriction region appears for a breakup but does not for non-breakup: the appearance of the constriction point is a sign of the breakup.
We further explore, in the third regime, analogy with critical phenomena in thermodynamic transitions [19, 20]. We show that the master curve, or scaling function for the breakup case is distinctively different from that for the non-breakup case with identifying three scaling exponents and find the exponents for the breakup case dependent on a length scale, which is a continuous number. This dependence is in contrast with (I) our previous studies [16, 17] and (II) the standard critical phenomena such as the ferromagnetic transition. This is because (I) in our previous studies the corresponding scaling exponents were found to be always the same value even if some length scales are changed and (II) in the standard critical phenomena the exponents usually depend on not continuous but discrete numbers, such as the number of components of the order parameter $n$, representing the symmetry, and dimensionality $d$. Since, in critical phenomena, such dependence is known to categorize universality classes, the dependence on a continuous number in the preset hydrodynamic analog indicates the existence of an uncountably infinite number of universality classes, as in some exotic cases known in critical phenomena. The deep analogy revealed in the present study suggests a promising avenue for the future study of self-similarity in PDEs by exploring many examples in natural phenomena and industrial processes, in which confined geometries are relevant.
@@
Figure 1: (a) Experimental setup. A metal disk of thickness $D_{0}$ (=1=1 mm) and radius $R$ (1010~to $12.5$ mm) falls in the cell of thickness $D$ (22 to 66 mm) filled with a viscous liquid of kinematic viscosity $\nu$ (100 to 1000 cS). The disk entrains air into the liquid, which finally detaches from the disk. The difference between $D$ and $D_{0}$ defines the liquid film thickness $e$. (b) Snapshots just before and at breakup illustrating the setting of axes, in the case with topology change for $(R,D_{0},e,\nu)=(10,3,0.5,100)$ in $m$ or cS.
@@
Figure 2: Snapshots of entrainment of air by a disk into liquid, leading to detachment of air from the disk for $R=10$ mm, $D_{0}=1$ $m$, $\nu=100$ cS. The liquid film thickness $e$ are 1, 1.5, and 2 mm, respectively, in (a) to (c). The time label 0 ms corresponds to $t=t_{d}$ defined in the text. In the three rightmost front-view photos reveal that a small bubble remains on the solid surface in (b) and (c) while such a bubble cannot be seen in (a): the breakup transition point lies between $e=$ 11 and 1.5 mm. Side-view shots near the detachment (0 ms) are separated by 6 ms, but are not synchronized with the front-view shots (a set of side-view snapshots are obtained from an experiment performed on a day different from the day on which the corresponding front-view snapshots but conducted for the same parameters).
@@ II.1 Experimental
In Fig. 1 (a), we show our experimental setup with explanation. The ranges of characteristic lengths, the radius and thickness of disk and the cell thickness, are as follows: R=1012.5R=10-12.5 mm, $D_{0}=1$ $m$, D=26D=2-6 mm. The cell width and height are much larger than the length scales $R,D_{0}$, and $D$ (typically 9 and 12 cm, respectively). We use polydimethylsiloxane (PDMS) for viscous liquid, where the range of kinematic viscosity $\nu=\eta/\rho$ is 100-1000 cS. The density $\rho$ and the surface tension $\gamma$ are slightly depending on viscosity $\eta$ ($\rho\simeq$ $0.97$ g/cm3 and $\gamma\simeq 20$ mN/m). The density $\rho_{s}$ of the metal disk (SUS430) is 7.7 g/cm3 with the density difference $\Delta\rho=\rho_{s}-\rho$. The cell is fabricated with acrylic plates of thickness 5 mm, using acrylic spacers whose thickness defines the cell thickness $D$.
@@
In the previous studies on the pre-detachment dynamics [16, 17], the sheet-forming with breakup and the corn-forming regime without breakup are typically observed for $D_{0}>\kappa^{-1}$ and $D_{0}<\kappa^{-1}$ at $e=0.5$ mm respectively, where $\kappa^{-1}=\sqrt{\gamma/(\rho g)}\simeq 1.8$ (In the study on the post-breakup dynamics [18], the case of the sheet-forming regime with breakup was studied). In the present study on the pre-detachment dynamics, we set $D_{0}$ to a fixed value 1 mm and change $e$ to find the transition from the breakup to the non-breakup in the corn-forming regime.
@@ II.2 Shape of the air-liquid interface $h(z,t)$
In Fig. 1 (b), we explain the setting of coordinates in the present study (we set $z_{c}=0$, so that $z-z_{c}=z$ and $z_{m}-z_{c}=z_{m}$). The air-liquid interface formed by air entrained by the disk is seen as the dark curve with a finite width. The width reflects the three-dimensional characters of the interface: the interface is concave or convex. The region where the width is thin, the interface is almost flat. We track the right and left inner edges of the thick curve and describe them by $x=h(z,t)$ and h(z,t)-h(z,t), respectively, except for the case of $e=0.5$ mm, for which we track instead the outer edges. We determine whether we track the inner or outer edge such that we can have a better collapse of shapes at different time after rescaling, which we discuss below. We call the minimum of the function $h(z,t)$ with respect to $z$ ”the constriction point” at which $(x,z)=(h_{m}(t),z_{m}(t))$, i.e., $h_{m}(t)=h(z_{m}(t),t)$.
II.3 Dynamics dependent on the film thickness $e$
In Fig. 2, we show snapshots of air-detachment near the air-solid contact for $e=1.0$, 1.5, and 2 mm, with the other parameters, $R$, $D_{0}$, and $\nu$ fixed (we have tried to analyze the case of $e=2.5$ mm, only to fail because the inner edge was not clear). In three rightmost front-view photos (labeled as 6 ms), detachment without topological change is observed for $e=1$ mm but that with topological change (i.e., breakup) for $e=1.5$ and $2$ mm: a small bubble remains on the solid surface in (b) and (c) while such a bubble cannot be seen in (a). In other words, ”the transition point between the breakup and non-breakup” or the breakup transition point lies between $e=$ 11 and 1.5 mm.
@@ II.3 Dynamics dependent on the film thickness $e$
In the previous studies [16] and [17], the cases of $D_{0}\gtrsim\kappa^{-1}$ and $D_{0}\lesssim\kappa^{-1}$ with $\kappa^{-1}=\sqrt{\gamma/(\rho g)}\simeq 1.8$ mm are distinguished by the terms sheet-forming and corn-forming detachments, and the former and latter cases are respectively regarded as the cases with and without topology transition and with and without appearance of the constriction point. However, in the present case, as seen in Fig. 2, both the cases with and without topology transition are observed for $D_{0}\lesssim\kappa^{-1}$ while the tip form seen from the side-view snapshots looks like a corn rather than a sheet and the constriction point, which is a sign of the breakup, can appear. In other words, we find that the transition from breakup to non-breakup can occur within the corn-forming detachment.
@@ II.4 Dynamics of characteristic length scales
In Fig. 3 (a1) to (a4), we show the dynamics of three characteristic length scales, $2h_{m}(t)$, $z_{m}(t)$, and $z_{G}(t)$, for different film thickness $e$. The quantity $z_{G}(t)$ is the position of the center of gravity of the disk measured from that at the critical time $t=t_{c}$, which slightly differs from $t_{d}$ $z_{G}(t_{c})=0$ by definition]. The critical time $t_{c}$ is determined by extrapolation as the time when $h_{m}$ becomes zero. The critical space-time point is represented as $(x,z,t)=(x_{c},z_{c},t_{c})$ with $x_{c}=h_{m}(t_{c})$ and $z_{c}=z_{m}(t_{c})$.
@@
[…]
We carefully selected the range of fitting indicated by a pair of cross marks for $h_{m}$ and $z_{m}$ to avoid including the data with large error bars and to use only the data with small error bars, by noting that (i) the error coming from the time resolution tends to become large near $t=t_{c}$ and (ii) the error tends to be large in $z_{m}$ compared with $h_{m}$. These issues are visualized by the error bars shown for the data at $e=1.0$ mm in (b2) and (b3): The error bars in the horizontal axis are based on the time resolution $\pm 0.5$ ms at 2000 fps; The error bars in the vertical axis are on the spacial resolution corresponding to 1 pixel of the images for $h_{m}$ but possibly corresponding to several pixels for $z_{m}$ because of the elongation explained below. In accordance with Remark (i) and (ii), we can confirm the following in (b2) and (b3): the error bars in the horizontal axis is visible near $t=t_{c}$, while the error bars in the vertical axis are not visible for $h_{m}$ but significant for small $z_{m}$. The reason for large errors in $z_{m}$ is as follows. The constriction region tends to elongate near the breakup and the thin thread of air thus formed detaches from the disk when $h_{m}$ is not zero (see Fig. 6 in Appendix A). Because of this elongation of the constriction region the determination of $z_{m}(t)$ tends to become difficult near breakup, which is reflected in larger standard deviation for $\Delta$ (compared with $\beta$) in Tab. 1.
As for the fitting of $z_{G}$, Fig. 3 (c) further justifies the linear fitting demonstrated in Fig. 3 (b1), supporting the following scaling law for $z_{G}(t)$ in agreement with the result established in a similar parameter range [17]:
@@ II.5 Self-similarity in the interface shape dynamics
In Fig. 4 (a1), we show the interfacial shape before rescaling at different times for $e=1$ mm. Interface shapes after rescaling are reasonably collapsed well on a master curve, especially near the constriction point, as seen in (a2). The collapse is significant for shapes before $t-t_{d}=-9$ ms as shown in (a3).
Figure 4: Shape change for $e=1$, 1.5, 2.0, and 0.5 mm for $(R,D_{0},\nu)=(10,1,100)$ in $m$ or cS. Since the right and left interfaces are the mirror image of the other, only the right interface after averaging is shown. The error bars in capturing the interface are less than the size of markers. The time development of interfacial shape before [after] rescaling is respectively shown in (a1) [(a2) and (a3)] for $e=1$ mm. We see a better collapse for times not too close to $t=t_{c}$ (before 9-9 ms) [see (a3)]. The corresponding plots for e=e=1.5, 2.0, and 0.5 mm are shown respectively in (b) to (d). The data labeled $R^{\ast}$ in (a3) are obtained for different parameter set $(R,D_{0},\nu)=(12.5,1,100)$, which collapse well on the master curve [the good collapse seems limited for the upper branch (of our focus), i.e., the shape above the constriction point].
Similar collapses are shown in (b) and (c) for $e=1.5$ and 2.0 mm, respectively. As $e$ increases the constriction region tends to elongate near breakup as announced, which can be confirmed by comparing (b1) with (c1) (see also Fig. 6 in Appendix A). This elongation is related to the above-mentioned breakup transition point: topology does change for $e=1.5$ and 2 mm with accompanied by breakup but it does not for $e=1.0$ mm. It seems that topological change (i.e., breakup) tends to elongate the constriction region near $t=t_{d}$. Because of this elongation, duration of a good collapse tends to be limited as $e$ increases [see (a2) to (c3) and (a3) to (c3)].
Reproducibility and universality are also demonstrated in Fig. 4. The data labeled $R^{\ast}$ in (a3) imply that the corresponding data are obtained for different $R$ with the other parameters fixed. In Fig. 4 (a3), even these data obtained for different $R$ collapse well on the master curve, which demonstrates the master curve is independent from (i.e., universal for change in) $R$ for a fixed $e$, indirectly representing a good reproducibility of the experiment.
In Fig. 4 (d1), we show the interfacial shape for $e=0.5$ mm, which is fundamentally different from cases of $e$ larger than 0.5 mm. In this case, as mentioned above, we need to track the outer edge instead of inner edge of the dark interface to observe collapse of the renormalized shape seen in (d2) [and more clearly in (d3)]. In addition, the constriction point does not appear [the point $(h_{m},$ $z_{m})$ is defined as the right contact point (at $y=0$) of the interface with the disk] but the tip forms a corn. However, the corn may not be axisymmetric but elliptic as reported in our previous study [17].
@@
Figure 5: Translated shape functions for $e=1.0,1.5,2.0$, and 0.5 mm on linear scales in the period of good collapse [(a1) to (d1)] and on log-log scales for 21-21 to 30-30 ms with a fitting line obtained by fitting the data in the region between the two black crosses for 21-21 ms [(a2) to (d2)]. For all cases, $(R,D_{0},\nu)=(10,1,100)$ in $m$ or cS.